# Neural network learns low-dimensional polynomials with SGD near the information-theoretic limit

### v1

@@ B.1.2 Verifying Assumption 3
    If $\mathrm{IE}(\sigma_{*})=1$. We require $\beta_{1}>0$ and $\sum_{j=1}^{q}j!\alpha_{j}\beta_{j}s^{j-1}>0$ for all $s>0$.
@@
    Else if $\mathrm{IE}((\sigma_{*})^{I})=1$ for some $2\leq I\leq C_{\sigma}$. We require $H(\sigma^{(I)}(\sigma^{(1)})^{I-1};0)$ is not 00 and has the same sign as $H((\sigma_{*})^{I};1)$. Also, $\sum_{j=2}^{q}j!\alpha_{j}\beta_{j}s^{j-1}>0$ for all $s>0$.
  • (III)
    Else if $\mathrm{IE}(\sigma_{*})=2$. We require $\beta_{2}>0$ and $\sum_{j=2}^{q}j!\alpha_{j}\beta_{j}s^{j-1}>0$ for all $s>0$.
  • (IV)
    Else $\mathrm{IE}((\sigma_{*})^{I})=2$ for some $2\leq I\leq C_{q}$. We require $H(\sigma^{(I)}(\sigma^{(1)})^{I-1};1)$ is not 00 and has the same sign as $H((\sigma_{*})^{I};2)$. Also, $\sum_{j=3}^{q}j!\alpha_{j}\beta_{j}s^{j-1}>0$ for all $s>0$.
Now we prove Lemma 3 which verifies the existence of an activation function that satisfies the assumptions above with non-zero probability. The construction does not depend on the link function itself, but only its degree $q$.
Proof of Lemma 3. Let $c$ be a sufficiently small constant, and $C_{\sigma}$ be the minimum odd integer with $C_{\sigma}\geq\max\{C_{q}+1,q+2,3\}$. With probability $1/2$, we choose the coefficients as $\beta_{1}\sim\mathrm{Unif}(\{\pm 1\})$, and $\beta_{j}\sim\mathrm{Unif}(\{-c,c\})$ for $2\leq j\leq C_{\sigma}$. Then, it is easy to see (I) and (III) are met with probability at least $2^{-q}$, because they hold when $\mathrm{sign}(\alpha_{j})=\mathrm{sign}(\beta_{j})$ holds in the summation.
By taking $c$ sufficiently small, we have
$H(\sigma^{(I)}(\sigma^{(1)})^{I-1};0)=\underbrace{i!\beta_{I}(\beta_{1})^{I-1}}_{\text{$\asymp c$}}+O(c).$ (B.6)
@@
When $I$ is even, by adjusting the sign of $\beta_{1}$, $H(\sigma^{(I)}(\sigma^{(1)})^{I-1};0)$ is not 00 and has the same sign as $H((\sigma_{*})^{I};1)$ with probability $\frac{1}{2}$. Note that the sign of $\beta_{1}$ is independent from whether $\sum_{j=2}^{q}j!\alpha_{j}\beta_{j}s^{j-1}>0$ for all $s>0$ holds. This holds with probability at least $2^{-q+1}$. Thus we verified (II) for even $I$.
In the same vein, we can verify (IV) for even $I\leq C_{q}\leq C_{\sigma}-1$. We have
$H(\sigma^{(I)}(\sigma^{(1)})^{I-1};1)=\underbrace{(I+1)!\beta_{I+1}(\beta_{1})^{I-1}}_{\text{$\asymp c$}}+O(c^{2}),$ (B.7)
and (IV) can be verified using the same argument.
Otherwise (also with probability $1/2$), we choose the coefficients as $\beta_{j}\sim\mathrm{Unif}(\{-c,c\})$ for $21\leq j\leq C_{\sigma}-$ 2 and $\beta_{C_{\sigma}-1}=\beta_{C_{\sigma}}=\pm 1$ to verify (II) and (IV) for odd $I$. It is easy to see that $\sum_{j=2}^{q}j!\alpha_{j}\beta_{j}s^{j-1}>0$ for all $s>0$ holds for (I) and $\sum_{j=3}^{q}j!\alpha_{j}\beta_{j}s^{j-1}>0$ for all $s>0$ for (III). In addition, $H((\mathsf{He}_{C_{\sigma}}+\mathsf{He}_{C_{\sigma}-1})^{(I)}((\mathsf{He}_{C_{\sigma}}+\mathsf{He}_{C_{\sigma}-1})^{(1)})^{I-1};0)>0$ and $H((\mathsf{He}_{C_{\sigma}}+\mathsf{He}_{C_{\sigma}-1})^{(I)}((\mathsf{He}_{C_{\sigma}}+\mathsf{He}_{C_{\sigma}-1})^{(1)})^{I-1};1)>0$. Therefore, by taking $c$ sufficiently small, flipping the sign of $\mathsf{He}_{C_{\sigma}}+\mathsf{He}_{C_{\sigma}-1}$ can change the sign of $H(\sigma^{(I)}(\sigma^{(1)})^{I-1};k)$ for both (II) and (IV) with odd $I$. Combining all cases yields the desired claim. ∎
B.2 Random Initialization

### v2

@@ B.1.1 For polynomial link functions
    If $I=1\Leftrightarrow\mathrm{IE}(\sigma_{*})=\mathrm{GE}(\sigma_{*})=p_{*}$.
    • Weak recovery: $\alpha_{p_{*}}\beta_{p_{*}}>0$ (covered by Assumption 3).
    • Strong recovery: $\sum_{j=p_{*}}^{q}j!\alpha_{j}\beta_{j}s^{j-1}>0$ for all $s>0$ (covered by Assumption 2).
    • Approximation (ridge regression): $\beta_{i}\neq 0$ for some $i\geq q$ (covered by Assumption 3).
@@
    If $2\leq I=\{\min\ i\mid\mathrm{IE}(\sigma_{*}^{I})=\mathrm{GE}(\sigma_{*})=p_{*}\}\leq C_{\sigma}$.
    • Weak recovery: $H((\sigma_{*})^{I};p_{*})H(\sigma^{(I)}(\sigma^{(1)})^{I-1};p_{*}-1)>0$ (covered by Assumption 3).
    • Strong recovery: $\sum_{j=p_{*}+1}^{q}j!\alpha_{j}\beta_{j}s^{j-1}>0$ for all $s>0$ (covered by Assumption 2).
    • Approximation: $\beta_{i}\neq 0$ for some $i\geq q$ (covered by Assumption 3).
Note that it is difficult to construct a deterministic activation function that satisfies all of the assumptions for any link function $\sigma_{*}$ (the simplest counterexample is to consider $-\sigma_{*}$ which flips the Hermite coefficients). Instead, we show the existence of randomized construction of such an activation function that satisfies all of the assumptions on the activation function simultaneously with constant probability, which entails that a subset of neurons can achieve strong recovery. The construction does not depend on properties of the link function itself except for its degree $q$.
Lemma 11.
There exists a randomized activation function sampled from a discrete set such that the above conditions hold with constant probability.
Proof.  Let $c$ be a sufficiently small constant only used in this proof and $C_{\sigma}$ be the minimum odd integer with $C_{\sigma}\geq\max\{C_{q}+1,q+2,3\}$, where $C_{q}$ was introduced in Proposition 6. With probability $\frac{1}{2}$, we let $\beta_{1}\sim\mathrm{Unif}(\{\pm 1\})$, and $\beta_{j}\sim\mathrm{Unif}(\{\pm c\})$ for $2\leq j\leq C_{\sigma}$. With probability $\frac{1}{2}$, we let $\beta_{j}\sim\mathrm{Unif}(\{\pm c\})$ for $21\leq j\leq C_{\sigma}-$ 2 and $\beta_{C_{\sigma}-1}=\beta_{C_{\sigma}}\sim\mathrm{Unif}(\{\pm 1\})$.
We first consider (I). When $\beta_{1}\sim\mathrm{Unif}(\{\pm 1\})$, and $\beta_{j}\sim\mathrm{Unif}(\{\pm c\})$ for $2\leq j\leq C_{\sigma}$, it is easy to see $\mathrm{sign}(\alpha_{j})=\mathrm{sign}(\beta_{j})$ for all $j=1,\cdots,q$ hold with probability at least $2^{-q}$, which is sufficient to satisfy (I).
We then consider (II). First focus on the case when $p_{*}=1$ and $I$ is even. When $\beta_{1}\sim\mathrm{Unif}(\{\pm 1\})$ and $\beta_{j}\sim\mathrm{Unif}(\{\pm c\})$ for $2\leq j\leq C_{\sigma}$, by taking $c$ sufficiently small, we have
$H(\sigma^{(I)}(\sigma^{(1)})^{I-1};0)=\underbrace{I!\beta_{I}(\beta_{1})^{I-1}}_{\text{$\asymp c$}}+O(c^{2}).$ (B.8)
@@
When $I$ is even, by adjusting the sign of $\beta_{1}$, $H(\sigma^{(I)}(\sigma^{(1)})^{I-1};0)$ is non-zero and has the same sign as $H((\sigma_{*})^{I};1)$ with probability $\frac{1}{2}$. Note that the sign of $\beta_{1}$ is independent from whether $\sum_{j=2}^{q}j!\alpha_{j}\beta_{j}s^{j-1}>0$ for all $s>0$ holds. This holds with probability at least $2^{-q+1}$. Thus we verified (II) for $p_{*}=1$ and even $I$.
For $p_{*}=1$ and odd $I$, consider $\beta_{j}\sim\mathrm{Unif}(\{\pm c\})$ for $21\leq j\leq C_{\sigma}-$ 2 and $\beta_{C_{\sigma}-1}=\beta_{C_{\sigma}}\sim\mathrm{Unif}(\{\pm 1\})$. Note that $\sum_{j=2}^{q}j!\alpha_{j}\beta_{j}s^{j-1}>0$ for all $s>0$ (this is the condition for strong recovery) and the condition for ridge regression also holds. Furthermore, the sign of $H((\mathsf{He}_{C_{\sigma}}+\mathsf{He}_{C_{\sigma}-1})^{(I)}((\mathsf{He}_{C_{\sigma}}+\mathsf{He}_{C_{\sigma}-1})^{(1)})^{I-1};0)$ is ±$\pm 1$ with equiprobability, independent of $\beta_{2},\dots,\beta_{q}$. Therefore, by taking $c$ sufficiently small, we can obtain the desired sign of $H(\sigma^{(I)}(\sigma^{(1)})^{I-1};0)$. Thus we proved (II) for $p_{*}=1$ and odd $I$.
Regarding (II) for $p_{*}=2$ and even $I$, when $\beta_{1}\sim\mathrm{Unif}(\{\pm 1\})$ and $\beta_{j}\sim\mathrm{Unif}(\{\pm c\})$ for $2\leq j\leq C_{\sigma}$, we have
$H(\sigma^{(I)}(\sigma^{(1)})^{I-1};1)=\underbrace{(I+1)!\beta_{I+1}(\beta_{1})^{I-1}}_{\text{$\asymp c$}}+O(c^{2}).$ (B.9)
Thus, similar to (II) with $p_{*}=1$ and even $I$, we get (II) for $p_{*}=2$ and even $I$.
Finally, consider (II) for $p_{*}=2$ and odd $I$. When $\beta_{j}\sim\mathrm{Unif}(\{\pm c\})$ for $21\leq j\leq C_{\sigma}-$ 2 and $\beta_{C_{\sigma}-1}=\beta_{C_{\sigma}}\sim\mathrm{Unif}(\{\pm 1\})$, the sign of $H((\mathsf{He}_{C_{\sigma}}+\mathsf{He}_{C_{\sigma}-1})^{(I)}((\mathsf{He}_{C_{\sigma}}+\mathsf{He}_{C_{\sigma}-1})^{(1)})^{I-1};1)$ is ±$\pm 1$ with equiprobability when $I$ is odd, and this term dominates the others in $H(\sigma^{(I)}(\sigma^{(1)})^{I-1};1)$. Thus, (II) for $p_{*}=2$ and odd $I$ holds similarly to (II) for $p_{*}=1$ and odd $I$.
Now we have obtained the assertion for all cases. ∎
B.1.2 For general link functions
Now we consider non-polynomial link functions with potentially large generative exponent $p_{*}=\mathrm{GE}(\sigma_{*})\geq 2$. For weak and strong recovery to succeed, the conditions on the activation function are essentially the same as those for polynomial link functions:
  • (I)
    If $I=1\Leftrightarrow\mathrm{IE}(\sigma_{*})=\mathrm{GE}(\sigma_{*})=p_{*}$.
    • Weak recovery: $\alpha_{p_{*}}\beta_{p_{*}}>0$.
    • Strong recovery: $\sum_{j=p_{*}}^{\infty}j!\alpha_{j}\beta_{j}s^{j-1}>0$ for all $s>0$,
  • (II)
    If $2\leq I=\{\min\ i\mid\mathrm{IE}(\sigma_{*}^{I})=\mathrm{GE}(\sigma_{*})=p_{*}\}\leq C_{\sigma}$.
    • Weak recovery: $H((\sigma_{*})^{I};p_{*})H(\sigma^{(I)}(\sigma^{(1)})^{I-1};p_{*}-1)>0$,
    • Strong recovery: $\sum_{j=p_{*}+1}^{\infty}j!\alpha_{j}\beta_{j}s^{j-1}>0$ for all $s>0$.
Due to the proof strategy (which uses Taylor expansion), we also require that all differentials and sum of expectations appearing in the following proofs are well-defined and bounded.
To approximate a non-polynomial $\sigma_{*}$, we introduce the following condition on the activation function. We sample $\sigma_{j}$ from a discrete set (with bounded cardinality). Let $J$ be an index set such that the coefficients of $\sigma_{j}\ (j\in J)$ satisfy the conditions above. Because we are selecting $\sigma_{j}$ from a discrete set, $|J|\simeq N$ holds. We introduce the following condition, which states that the target single-index model can be well-approximated by a linear combination of student neurons.
Assumption 4.
When $b_{j}\sim\mathrm{Unif}([-C_{b},C_{b}])$ where $(C_{b}=\mathrm{polylog}(d))$ and $\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{T_{2}}\sim\mathcal{N}(0,\boldsymbol{I}_{d})$, there exists a set of coefficients $a_{1},\dots,a_{|J|}$ such that
$\frac{1}{T_{2}}\sum_{i=1}^{T_{2}}\bigg{(}\frac{1}{|J|}\sum_{j\in J}a_{j}\sigma_{j}(\boldsymbol{\theta}_{j}^{\top}\boldsymbol{x}_{i}+b_{j})-\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}_{i})\bigg{)}^{2}\lesssim\varepsilon^{2},$ (B.10)
holds with coefficients of reasonable magnitudes $\sum_{j\in J}a_{j}^{2}=\Theta(|J|)$ with high probability (w.r.t. the randomness of $b_{j}$ and $\boldsymbol{x}_{i}$). Moreover, $\mathbb{E}_{\boldsymbol{x}}[\sigma_{j}(\boldsymbol{\theta}_{j}^{\top}\boldsymbol{x}+b_{j})^{4}]\leq\mathrm{polylog}(d)$ for all $j$ with high probability (w.r.t. the randomness of $b_{j}$).
The following lemma states that we can design a randomized activation function that satisfies all of the above assumptions with probability $\Omega(1)$, as long as the link function $\sigma$ satisfies Assumption 4 for $\sigma=\mathrm{ReLU}$. In other words, we are able to cover the class of link functions $\sigma$ that can be efficiently approximated by a two-layer ReLU network. Since the general link functions are not included in a compact space, we do not have an upper bound of exponent to obtain $\mathrm{IE}(\sigma_{*}^{I})=\mathrm{GE}(\sigma_{*})$ as we had $C_{q}$ in the polynomial case. Consequently, our student activation is not entirely agnostic to the link function $\sigma_{*}$, as we require knowledge of $p$ (information exponent), $p_{*}$ and $I$.
Lemma 12.
Suppose the target link function $\sigma_{*}$ satisfies Assumption 4 for $\sigma_{j}=\mathrm{ReLU}$. There exists a randomized activation sampled from a discrete set such that the above conditions hold with constant probability.
Before we sketch the design of activation function, we present the following approximation result from [BBSS22], which establishes that Assumption 4 with $\sigma_{j}=\mathrm{ReLU}$ is satisfied for broad class of functions, according to Lemma 4.4 and 4.5 of [BBSS22]. Specifically, taking $\tau=1/2$ and $\lambda=N^{-1}$ yields that $\mathbb{E}_{\boldsymbol{x}}[(\frac{1}{|J|}\sum_{j\in J}a_{j}\sigma_{j}(\boldsymbol{\theta}^{\top}\boldsymbol{x}+b_{j})-\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}))^{2}]\leq N^{-\frac{2}{7}}$. Although they sample $b_{j}$ from Gaussian $\mathcal{N}(0,2)$, the result translates to uniform sampling of biases from $[-C_{b},C_{b}]$ by introducing additional logarithmic factor.
Lemma 13 (Lemma 4.4, 4.5 of [BBSS22]).
Suppose that $\mathbb{E}_{z\sim\mathcal{N}(0,1)}[\sigma_{*}(z)^{4}]$, $\mathbb{E}_{z\sim\mathcal{N}(0,1)}[\sigma_{*}^{\prime\prime}(z)^{4}]<\infty$. Then, Assumption 4 with $\sigma_{j}=\mathrm{ReLU}$ holds with $\varepsilon=N^{-\frac{1}{7}}$ and $C_{b}\simeq\sqrt{\log d}$.
Proof of Lemma 12. We show the existence of suitable $\sigma$ in two steps: first we construct a randomized polynomial activation function that satisfies conditions (I)(II) with constant probability; then we add a small ReLU perturbation so that the activation can approximate non-polynomial $\sigma_{*}$.
Recall $p\in\mathbb{N}_{+}$ is the information exponent of $\sigma_{*}$. We first show that there exists a randomized polynomial activation that satisfies the conditions for weak and strong recovery with probability $\Omega(1)$. Note that the issue of differentiability and bounded moment is avoided when we focus on the polynomial activation functions. We specify the following two distributions. With probability $\frac{1}{2}$, let $\beta_{1}\sim\mathrm{Unif}(\{-1,1\})$, $\beta_{j}\sim\mathrm{Unif}(\{-c,c\})$ for $j=1,\cdots,p_{*}+I-1$ and $\beta_{j}=0$ otherwise, where $c>0$ is a sufficiently small constant. With probability $\frac{1}{2}$, let $\beta_{1}=\mathrm{Unif}(\{-1,1\})$, $\beta_{2}=\mathrm{Unif}(\{-c,c\})$, $\beta_{j}=\mathrm{Unif}(\{-c^{2},c^{2}\})$ for all $2\leq j\leq(p_{*}+I)\lor p$ for a sufficiently small constant $c>0$, and $\beta_{j}=0$ otherwise.
Regarding (I), consider the case when the coefficients are sampled from the first distribution, and $|\beta_{j}|\ll 1$ except for $j=p_{*}$. Then, $\sum_{j=p_{*}}^{\infty}j!\alpha_{j}\beta_{j}s^{j-1}\approx p_{*}!\alpha_{p_{*}}\beta_{p_{*}}s^{p_{*}}$. Choosing the sign of $\beta_{p_{*}}$, we have that the assumption holds with probability $\Omega(1)$.
Regarding (II) with even $I$, consider coefficients sampled from the first distribution, and $\mathrm{Sign}(\beta_{j})=\mathrm{Sign}(\alpha_{j})$ for $j\leq(p_{*}+I-1)\lor p$. Then, $\sum_{j=p_{*}+1}^{\infty}j!\alpha_{j}\beta_{j}s^{j-1}>0$ for all $s>0$. Also, similarly to (B.8),
$H(\sigma^{(I)}(\sigma^{(1)})^{I-1};p_{*}-1)=\underbrace{i!\beta_{I+p_{*}-1}(\beta_{1})^{I-1}}_{\text{$\asymp c$}}+O(c^{2}).$ (B.11)
By flipping the sign of $\beta_{1}$, we can change the sign of $H(\sigma^{(I)}(\sigma^{(1)})^{I-1};p_{*}-1)$. Thus, (II) for even $I$ is satisfied by a randomized choice of $\beta_{1}$.
For (II) with odd $I$, consider coefficients sampled from the second distribution, and $\mathrm{Sign}(\beta_{j})=\mathrm{Sign}(\alpha_{j})$ for $j\leq(p_{*}+I)\lor p$. Then, $\sum_{j=p_{*}+1}^{\infty}j!\alpha_{j}\beta_{j}s^{j-1}>0$ for all $s>0$.
$H(\sigma^{(I)}(\sigma^{(1)})^{I-1};p_{*}-1)=\frac{1}{(p_{*}-1)!}\mathbb{E}[\sigma^{(I)}(\sigma^{(1)})^{I-1}\mathsf{He}_{p_{*}-1}]$ (B.12)
$=\frac{1}{(p_{*}-1)!}\mathbb{E}[(I-1)(\beta_{p_{*}+I}\mathsf{He}_{p_{*}+I})^{(I)}(\beta_{2}\mathsf{He}_{2})^{(1)}(\beta_{1})^{I-2}\mathsf{He}_{p_{*}-1}]+O(c^{4}).$ (B.13)
$=\underbrace{\frac{2(I-1)\beta_{p_{*}+I}\beta_{2}(\beta_{1})^{I-2}(p_{*}+I)!}{(p_{*}-1)!}}_{\text{$\asymp c^{3}$}}+O(c^{4}).$ (B.14)
By flipping the sign of $\beta_{1}$, we can change the sign of $H(\sigma^{(I)}(\sigma^{(1)})^{I-1};p_{*}-1)$. Thus, (II) for odd $I$ is satisfied by a randomized choice of $\beta_{1}$.
Therefore, we have constructed a randomized polynomial activation $\sigma$ that satisfies all of the conditions for weak and strong recovery. Now we provide a sketch of reasoning that when the link function $\sigma_{*}$ is well-approximated by ReLU as specified in Assumption 4, we can find some $\sigma$ that additionally satisfies Assumption 4 by introducing a small ReLU component. Specifically, we add $c_{\mathrm{R}}\cdot\mathrm{ReLU}$ to the activation function with probability $\frac{1}{2}$, with a sufficiently small $c_{\mathrm{R}}=\tilde{\Omega}(1)$, e.g., $c_{\mathrm{R}}=(\text{log}d)^{-C}$ for some $C>0$. When a two-layer ReLU network approximates $\sigma_{*}$ that satisfies Assumption 4, by using the neurons with added ReLU component, $\sigma_{*}$ can be approximated up to some polynomial residual with degree $(p_{*}+I)\lor p$. And by using the remaining polynomial neurons, we can approximate the additional polynomial terms in $\sigma_{*}$ (see Lemma 22,23). Subtracting the latter from the former, we obtain the desired approximation result. When $c_{\mathrm{R}}$ is sufficiently small, this additional term does not impact the conditions for weak and strong recovery and the moment calculations; similarly, since $c_{\mathrm{R}}\ll 1$ we may discard this non-smooth term before Taylor expansion without affecting the analysis of optimization dynamics. We remark that to avoid such unnatural design of activation function, we can also train the first-layer parameters using a polynomial activation specified above, and then perturb it before the second-layer training to enhance the approximation ability — such strategy has also been employed in prior layer-wise training analysis [AAM22]. ∎
B.1.3 More Discussion on Assumption 2
Assumption 2 requires $H(\sigma^{(I)}(\sigma^{(1)})^{I-1};p_{*}-1)$ is not zero and has the same sign as $H(\sigma_{*}^{I};p_{*})$. We remark that if we allow a negative momentum parameter larger than $1$, i.e., setting $\xi^{2(t+1)}=1+c_{\xi}d^{-\frac{(p_{*}-2)_{+}}{2}}$, we can negate the opposite sign of $H(\sigma^{(I)}(\sigma^{(1)})^{I-1};p_{*}-1)$ (see Lemma 16), and the subsequent analysis still holds. Therefore, what we essentially need is $H(\sigma^{(i)}(\sigma^{(1)})^{i-1};k)\neq 0$. Lemma 3 confirms that it is satisfied by almost all polynomials:
Proof of Lemma 3. We note that $H(\sigma^{(i)}(\sigma^{(1)})^{i-1};k)=\mathbb{E}[\sigma^{(i)}(\sigma^{(1)})^{i-1}\mathsf{He}_{k}]$ is a polynomial of $\{\beta_{j}\}_{j=0}^{C_{\sigma}}$. This polynomial is not identically equal to zero. To confirm this, consider $\sigma=x^{C_{\sigma}}+x^{C_{\sigma}-1}$. Because $\sigma^{(i)}(\sigma^{(1)})^{i-1}$ is expanded as a sum of $x^{l}(i(C_{\sigma}-3)\leq l\leq i(C_{\sigma}-2)+1$ with positive coefficients and each $x^{l}$ is a sum of $\mathsf{He}_{l},\mathsf{He}_{l-2}\cdots$ with positive coefficients, $\sigma^{(i)}(\sigma^{(1)})^{i-1}$ has all positive Hermite coefficients for degree $10,1,\cdots,i(C_{\sigma}-2)+$ 1. If $k\leq i(C_{\sigma}-2)+1$, this choice of $\sigma$ yields $H(\sigma^{(i)}(\sigma^{(1)})^{i-1};k)>0$, which confirms that $H(\sigma^{(i)}(\sigma^{(1)})^{i-1};k)$ as a polynomial of $\{\beta_{j}\}_{j=0}^{C_{\sigma}}$ is not identically equal to zero. Hence the assertion follows from so-called Schwartz–Zippel Lemma [Sch80], or the fact that zeros of a non-zero polynomial form a measure-zero set. ∎
B.2 Initialization